# Beyond Browsing: API-Based Web Agents

### v2

@@ Abstract
Yueqi Song, Frank Xu, Shuyan Zhou, Graham Neubig {yueqis,gneubig}@cs.cmu.edu Carnegie Mellon University
@@ 1 Introduction
At the same time, not all websites have extensive API support, in which case web browsing actions may still be required. To address these cases, we explore a hybrid approach that combines API-Based Agents with Browsing Agents, as described in Figure 1. By implementing an agent capable of interleaving API calls and web browsing, we found that agents benefit from the flexibility of this hybrid model. When APIs are available and well-documented, the agent can directly interact with the web services. For websites with limited API support, the agent seamlessly interleaves API calling and browsing to ensure task completion.
We evaluated our API-Based and Hybrid Agents on WebArena, a benchmark for real-world web tasks (Zhou et al., 2024), and the results are shown in Figure 1. Our experiments revealed three key findings: (1) The API-Based Agent outperforms the Browsing Agent on WebArena by around 15% on average. (2) The API-Based Agent yields a higher success rate on websites with good API support (e.g., Gitlab) compared to those with limited API support (e.g., Reddit). This result underscores the importance of developing comprehensive API support for more accurate and efficient web task automation in the future. (3) The Hybrid Agent outperforms solely Browsing and solely API-Based Agents, further improving accuracy by 6% compared to the API-Based Agent. By dynamically interleaving approaches, the Hybrid Agent is able to provide more consistent and reliable outcomes.
In sum, our results suggest that allowing agents to interact with APIs, interfaces designed specifically for machines, is often preferable or at least complementary to direct interaction with graphical interfaces designed for humans.
@@ 5.2 API Statistics for WebArena Sites
Table 1: Number of endpoints, and quality of API and documentation for WebArena websites.
@@ 5.4 Evaluation Framework
Table 2: Agents’ performances across WebArena tasks. Note that SteP uses prompts inspired specifically by WebArena tasks, while other agents are task-agnostic. We achieve the highest accuracy among task-agnostic agents.
@@ 6.1 Main Results
The Hybrid Agent, integrating both API calling and web browsing, outperformed the other agents on many websites. The agent’s ability to interleave API calling and web browsing proved beneficial. API calling delivers high performance for web tasks when well-supported APIs are available, while web browsing serves as a backup when API endpoints are unavailable or incomplete. Even if the website provides comprehensive APIs, there might be corner cases where APIs are not supportive. Thus, relying on web browsing is still needed for tasks that would otherwise fail through API-only interactions. Table 3 documents the frequency of each type actions of the Hybrid Agent: it chooses to do both Browsing and API in 77.7% of WebArena tasks, and it shows higher accuracy when choosing API only and API+browsing. More detailed analysis on action types, steps and cost and case studies are in Appendix A.4 and A.5.
@@
Browsing only 14.3 21.6
API only 8.0 38.5
Browsing+API 77.7 38.2
@@ 6.2 Does API Quality Matter?
Conversely, low-quality APIs, characterized by incomplete functionality or ambiguous documentation, can significantly degrade performance. In such cases, the absence of necessary endpoints may prevent the API-Based Agent from completing tasks and force the Hybrid Agent to resort to browsing. Moreover, poorly documented APIs can result in misusing parameters and headers, further reducing the effectiveness of the agent. This highlights the importance for websites to maintain comprehensive and well-documented API support.
@@
Moreover, API quality can also correlate with the performance of Browsing Agents. This may be because websites with well-implemented APIs often have clean, user-friendly interfaces, which benefit machine agents when interacting with the web interface. Good API practices suggest a thoughtful design process that tends to carry over into the overall user interface and experience, allowing the Browsing Agent to more easily parse and interact with the website’s elements. As a result, both API-Based and Browsing Agents are able to function more effectively in environments where high API standards are maintained.
7 Conclusion and Future Work
In this paper, we propose new web agents that use APIs instead of traditional browsers. We find that API-Based Agents outperform Browsing Agents, especially on sites with good API support. Thus we further propose an agent capable of interleaving API calling and browsing that empirically outperforms agents that only use one of the two interfaces.
For future work, we aim to explore automatically inducing APIs using methods such as Agent Workflow Memory (AWM) (Wang et al., 2024e). These methods could identify and generate API calls for websites lacking formal API support, further expanding the applicability and efficiency of API-Based Agents. By automating the discovery and utilization of APIs, we envision even more robust agents capable of handling diverse web tasks without reliance on interaction through browsing.
@@ 8 Limitations
Evaluation Benchmark
In this paper, we evaluate web agents exclusively on WebArena tasks. While WebArena offers realistic and diverse challenges, the number and variety of tasks may be limited. Other benchmarks, such as Webshop (Yao et al., 2022), MiniWoB (Shi et al., 2017), Mind2Web (Deng et al., 2023), WebVoyager (He et al., 2024b), and VisualWebArena (Koh et al., 2024a), provide alternative valuable evaluation platforms. However, as discussed in Section 2.1, WebArena aligns more closely with real-world scenarios and our use case, while other benchmarks lack support for API calling. For example, VisualWebArena is less applicable to our study because WebArena APIs lack support for interacting with images, a core component of VisualWebArena tasks.
@@
Unlike Browsing Agents, which can adapt to new websites without manual intervention, the API-Based Agent requires additional effort to integrate the necessary APIs documentation to the action space of the agent for each website. This manual integration process increases complexity, particularly when the agent must support a wide range of websites, limiting scalability compared to agents that rely solely on web browsing for interactions. However, future advancements in automated API scraping and documentation generation could eliminate this bottleneck, allowing for more scalable and flexible API-Based Agents.
@@ A.4 Additional Analysis
Figure 6: Error analysis on 100 randomly sampled WebArena tasks.
Error Analysis
We randomly sampled 100 tasks from the WebArena benchmark and performed error analysis on the API-Based Agent. Figure 6 shows the distribution of error categories among these 100 tasks. We found that 33% of the tasks are correctly performed with only API calling, 50% are unsolvable with solely APIs, 6% are incorrect due to incorrect task understanding, and 11% are incorrect due to error in calling APIs such as mal-formatting and wrong input. In other words, among the 50 API solvable tasks, 66% are performed correctly by the API-Based Agent. This showcases the strong capability of the API-Based Agent when given sufficient APIs to solve the task.
Additionally, the average API calls required to solve the API solvable tasks are 2.1 API calls, demonstrating how API calling could reduce operational complexity for web tasks. Although the average number of steps the API-Based Agent took to complete WebArena tasks is 7.8 steps, most of the steps were taken by the agent to retrieve API documentation, resolve errors from it’s previous generations, or verify it’s outputs.
A.5 Case Studies
In this section, we analyze two contrasting instances as shown in Figure 7 and Figure 8, where the Hybrid Agent and API-Based Agent exhibited different levels of performance on WebArena tasks. These case studies highlight the strengths and weaknesses of each agent, demonstrating scenarios where hybrid browsing outperforms API-only or browsing-only approaches, as well as cases where the API-Based Agent excels over the hybrid method.
Case 1
Figure 8: Case 2: the API-Based Agent succeeds while the Browsing Agent and the Hybrid Agent fails.
Case 2
A.6 API-Based Agent Prompt
@@
A.7 Hybrid Agent Prompt

### v3

@@ Abstract
Yueqi Song, Frank Xu, Shuyan Zhou, Graham Neubig {yueqis,gneubig}@cs.cmu.edu Carnegie Mellon University https://yueqis.github.io/API-Based-Agent/
@@ 1 Introduction
However, not all websites have extensive API support, in which case web browsing actions may still be required. To overcome this limitation, we propose a hybrid approach that combines API-Based Agents with Browsing Agents, as depicted in Figure 1. Rather than choosing between API calling and web browsing at the task level, our Hybrid Agent is capable of dynamically interleaving both actions. We found that agents benefit from the flexibility of this hybrid model. When APIs are available and well-documented, the agent can directly interact with the web services. For websites with limited API support, the agent seamlessly interleaves API calling and browsing, combining the power of both to complete each task.
We evaluated our API-Based and Hybrid Agents on WebArena, a benchmark for real-world web tasks (Zhou et al., 2024), and the results are shown in Figure 1. Our experiments revealed three key findings: (1) The API-Based Agent outperforms the Browsing Agent on WebArena by around 15% on average. (2) The API-Based Agent yields a higher success rate on websites with good API support (e.g., Gitlab) compared to those with limited API support (e.g., Reddit). This result underscores the importance of developing comprehensive API support for more accurate and efficient web task automation in the future. (3) The Hybrid Agent outperforms solely Browsing and solely API-Based Agents, further improving accuracy by 10% compared to the API-Based Agent. By dynamically interleaving approaches, the Hybrid Agent is able to provide more consistent and reliable outcomes.
In sum, our results suggest that allowing agents to interact with APIs, interfaces designed specifically for machines is often preferable or at least complementary to direct interaction with graphical interfaces designed for humans.
@@ 5.2 API Statistics for WebArena Sites
Table 1: Number of endpoints, and quality of API and documentation for WebArena websites. See Section 6.2 for more discussions.
@@ 5.4 Evaluation Framework
Table 2: Agents’ performances on WebArena. Note that SteP uses prompts inspired specifically by WebArena tasks, while other agents are task-agnostic. Our Hybrid Agent achieve the highest accuracy among these agents. We report these results using a set of APIs introduced by the authors to mimic the official Reddit website, constructed task-agnostically without access to WebArena tasks. See Section 6.2 for more discussions.
@@ 6.1 Main Results
The Hybrid Agent, integrating both API calling and web browsing, outperformed the Browsing and API-Based Agents on all websites. The agent’s ability to interleave API calling and web browsing proved beneficial. API calling delivers high performance for web tasks when well-supported APIs are available, while web browsing serves as a backup when API endpoints are unavailable or incomplete. Even if the website provides comprehensive APIs, there might be corner cases where APIs are not supportive. Thus, relying on web browsing is still needed for tasks that would otherwise fail through API-only interactions. Table 3 show the frequency of each action type of the Hybrid Agent: it chooses to do both Browsing and API in 77.7% of WebArena tasks, and it shows higher accuracy when choosing API only and API+browsing. More detailed analysis on action types, steps and cost and case studies are in Appendix A.4 and 6.4.
@@
Browsing only 14.3 25.0
API only 8.0 40.0
Browsing+API 77.7 42.0
@@ 6.2 Does API Quality Matter?
Conversely, low-quality APIs, characterized by incomplete functionality or ambiguous documentation, can significantly degrade performance. In such cases, the absence of necessary endpoints may prevent the API-Based Agent from completing tasks. Moreover, poorly documented APIs can result in misusing parameters and headers, further reducing the effectiveness of the agent. This highlights the importance for websites to maintain comprehensive and well-documented API support.
@@
Figure 4: Error analysis on 100 WebArena tasks.
6.3 Error Analysis
We randomly sampled 100 tasks from WebArena and performed error analysis on the API-Based Agent. Figure 4 shows the distribution of error categories among these 100 tasks. We found that 33% of the tasks are correctly performed with only API calling, 50% are unsolvable with solely APIs, 6% are incorrect due to incorrect task understanding, and 11% are incorrect due to error in calling APIs such as mal-formatting and wrong input. In other words, among the 50 API solvable tasks, 66% are performed correctly by the API-Based Agent. This showcases the strong capability of the API-Based Agent when given sufficient APIs to solve the task.
Additionally, the average API calls required to solve API solvable tasks are 2.1 API calls, demonstrating how API calling could reduce operational complexity for web tasks. Although the API-Based Agent took an average of 7.8 steps to complete WebArena tasks, most of the steps were taken to retrieve API documentation, resolve errors from it’s previous generations, or verify it’s outputs.
6.4 Case Studies
We analyze two contrasting instances as shown in Figure 5 and Figure 6, where the Hybrid Agent and API-Based Agent exhibited different levels of performance on WebArena tasks. These case studies highlight the strengths and weaknesses of each agent, demonstrating scenarios where hybrid browsing outperforms API-only or browsing-only approaches, as well as cases where the API-Based Agent excels over the hybrid method.
Case 1
Figure 6: The API-Based Agent succeeds while the Browsing Agent and the Hybrid Agent fails.
Case 2
7 Conclusion
In this paper, we propose new web agents that use APIs instead of traditional browsers. We find that API-Based Agents outperform Browsing Agents, especially on websites with good API support. Thus we further propose Hybrid Agents capable of interleaving API calling and browsing that empirically outperforms agents that only use one of the two interfaces.
@@ 8 Limitations
Unlike Browsing Agents, which can adapt to new websites without manual intervention, the API-Based Agent requires additional effort to integrate the necessary APIs documentation to the action space of the agent for each website. This manual integration process increases complexity, particularly when the agent must support a wide range of websites, limiting scalability compared to agents that rely solely on web browsing for interactions. However, future advancements could explore automatically inducing APIs using methods such as Agent Workflow Memory (AWM) (Wang et al., 2024e) and self-improving (Zheng et al., 2025). These methods could identify and generate API calls for websites lacking formal API support, further expanding the applicability and efficiency of API-Based Agents. By automating the discovery and utilization of APIs, we envision even more robust agents capable of handling diverse web tasks without reliance on interaction through browsing.
Evaluation Benchmark
In this paper, we evaluate web agents exclusively on WebArena tasks. While WebArena offers realistic and diverse challenges, the number and variety of tasks may be limited. Other benchmarks, such as Webshop (Yao et al., 2022), MiniWoB (Shi et al., 2017), Mind2Web (Deng et al., 2023), WebVoyager (He et al., 2024b), and VisualWebArena (Koh et al., 2024a), provide alternative valuable evaluation platforms. However, as discussed in Section 2.1, WebArena aligns more closely with real-world scenarios and our use case, while other benchmarks lack support for API calling. For example, VisualWebArena is less applicable to our study because WebArena APIs lack support for interacting with images, a core component of VisualWebArena tasks. This could be potentially solved by aforementioned future approaches to automatically induce APIs that support image interactions. Nevertheless, we consider our work as a foundation for such future explorations.
@@ A.3 Obtaining APIs of WebArena Websites
  • Multi-Website Tasks: we provide APIs from all websites included in the task to the agents, where we explicitly state which set of APIs belongs to which website. This could allow agents to identify the correct set of APIs to use when transitioning between websites. It is also worth noting that the framework of our agents supports a unified workspace that allows the agents to carry over the information from one website to another.
@@ A.4 Additional Analysis
A.5 API-Based Agent Prompt
@@
A.6 Hybrid Agent Prompt